Solve using the square root property:
(x + 3)^2 = 36
please i need a good example to follow, for my other problems.

Found 2 solutions by checkley75, jim_thompson5910:
Answer by checkley75(3666)

(X+3)^2=36
(X+3)(X+3)=36
X^2+6X+9=36
X^2+6X+9-36=0
X^2+6X-27=0
(X+9)(X-3)=0
X+9=0
X=-9 ANSWER.
X-3=0
X=3 ANSWER.

Answer by jim_thompson5910(35256)

Here's another way to do it:

take the square root of both sides

So x+3 equals positive 6 or negative 6 so:
Evaluate x+3=6
Subtract 3 from both sides
Evaluate the right side
So x=3
Evaluate x+3=-6
Subtract 3 from both sides
Evaluate the right side
So our answer is:
x=3 or x=-9